EU Server node latency (Resolved)

Certain VM's and shared hosting are affected by this outage.
We are working to correct the problem.

Thank you for your patience.

Effects - This issue does not effect your relevant products/services
Affecting Server - Vps Control Panel (CA)
Priority - Critical
Date - 03/27/2017 18:40 - 04/08/2017 06:56
Last Updated - 03/27/2017 18:41

Server down (Resolved)

We are having troubles with this server and all hosted site on this server are down.
We are working the issue to resolve.

Server is now temporarily up.
Planning to move all hosted sites to new server.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Medium
Date - 01/14/2017 10:46 - 03/27/2017 18:40
Last Updated - 01/14/2017 12:08

Host Node slave03fr.bibihost.com Down (Resolved)

Host Node slave03fr.bibihost.com is Down and we are investigating.

This affects all VPS's on that node.

Effects - This issue does not effect your relevant products/services
Affecting System - slave03fr.bibihost.com failure
Priority - Critical
Date - 01/14/2017 09:13 - 03/27/2017 18:39
Last Updated - 01/14/2017 09:15

Outage with subnet 95.141.39.xxx (Resolved)

Discrepancy with IP address block caused issue.
We have taken steps to ensure this doesn't happen again.
All services all up.
Thank you for your patience!

Reported outage with subnet 95.141.39.xxx

Effects - This issue does not effect your relevant products/services
Affecting Other - Outage with subnet 95.141.39.xxx
Priority - Critical
Date - 12/01/2016 01:07 - 12/01/2016 13:14
Last Updated - 12/01/2016 06:20

Connectivity problems (Resolved)

Discrepancy with IP address block caused issue.
We have taken steps to ensure this doesn't happen again.
All services all up.
Thank you for your patience!

Looks like we are still working through some connectivity problems with cp245 and cp246
Seeking a solution...

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 11/30/2016 16:50 - 12/01/2016 13:16
Last Updated - 12/01/2016 06:20

Connectivity problems (Resolved)

Discrepancy with IP address block caused issue.
We have taken steps to ensure this doesn't happen again.
All services all up.
Thank you for your patience!

Looks like we are still working through some connectivity problems with cp245 and cp246
Seeking a solution...

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 11/30/2016 16:48 - 12/01/2016 13:15
Last Updated - 12/01/2016 06:19

Host node bad hard drive in array (Resolved)

Bad drive has been replaced and data rebuilt.
We are currently monitoring the system to ensure stability.



A bad hard drive was found on the host node server affecting shared hosting on cp245 and cp246 as well as VM's on Milano01.
We are working to replace the bad drive and rebuild data.
We will notify when work is complete.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 11/29/2016 13:38 - 11/30/2016 00:00
Last Updated - 11/30/2016 05:06

Host node bad hard drive in array (Resolved)

Bad drive has been replaced and data rebuilt.
We are currently monitoring the system to ensure stability.


A bad hard drive was found on the host node server affecting shared hosting on cp245 and cp246 as well as VM's on Milano01.
We are working to replace the bad drive and rebuild data.
We will notify when work is complete.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 11/29/2016 13:44 - 11/30/2016 00:00
Last Updated - 11/30/2016 05:05

Critical Security Updates to servers (Resolved)

UPDATE:

All patches applied. All servers rebooted. All servers and VM's back up





We need to perform various maintenance to each of these servers which is going to require a reboot. This is to patch some critical security vulnerabilities, while we have applied the workarounds/patches these are not 100% effective and only work against known PoC's.

We believe this is going to take around 30 minutes per server and are allocating the following 2 hour window in which the servers will be rebooted.

Start: 11th Nov 2016 02:00 UTC
End: 11th Nov 2016 04:00 UTC

Effects - This issue does not effect your relevant products/services
Affecting Other - ca01, milano01, slave01us, slave03fr and slave04fr
Priority - High
Date - 11/11/2016 02:00 - 11/11/2016 04:00
Last Updated - 11/11/2016 03:26

bgp flapping due to massive DDoS attack (Resolved)


We noticed bgp flapping due to massive DDoS incoming on all internal Service.
We mitigated the attack.
Only the website is not reachable from some location.
All other service are restored.





All issues have been resolved.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 09/22/2016 14:25
Last Updated - 09/22/2016 14:35

All servers upgraded apache and php then rebo (Resolved)

All cPanel servers upgraded apache and php then rebooted

Effects - This issue does not effect your relevant products/services
Affecting Other - All cPanel servers
Priority - Low
Date - 10/30/2015 06:07 - 10/30/2015 06:40
Last Updated - 10/30/2015 06:41

Italy DC Down (Resolved)

The Italian data center is experiencing issues with connectivity at the moment.

Update:

We're receiving huge DDoS on most our ip block, we're working to mitigate it

Update:

We have mitigated almost all of the attack and services should be running again.

Effects - This issue does not effect your relevant products/services
Affecting Other - Italy DC Down
Priority - Critical
Date - 10/14/2015 13:06
Last Updated - 10/14/2015 16:35

Server down (Resolved)

Data center has null routed this servers IP due to a few customers abusing their websites.

All websites have been moved to new servers and you will be notified soon as to your new location.

We are working diligently to completely resolve this problem.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 08/04/2015 11:44 - 08/05/2015 12:34
Last Updated - 08/04/2015 11:46

mySQL server down (Resolved)

We are currently working to resolve this problem and will update this area when resolved.

Problem has been resolved.
Thank you for your patience.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 08/28/2014 07:31 - 08/28/2014 07:56
Last Updated - 08/28/2014 07:56

Server not responding (Resolved)

We are aware of the problem and working to correct it.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 06/28/2014 13:13 - 06/28/2014 13:20
Last Updated - 06/28/2014 13:20

Node issue (Resolved)

We're working on it

------------------------------------

It was a filesystem issue and should be now resolved.

------------------------------------

A full fix will require a proper fsck to be run as we still see some errors logged. We'll try to do it by night since it'll require the node's partitions to be unmounted and obviously everybody on that node to go down.

Effects - This issue does not effect your relevant products/services
Affecting System - Milano01
Priority - Critical
Date - 03/25/2013 10:20 - 04/16/2013 14:33
Last Updated - 03/25/2013 14:26

Planned maintenance (Resolved)

A ram upgrade is scheduled on this node. Because of the necessary reboot, there will be some minutes of downtime.


-----------------------------------------------------------------------------------

There's an unexpected issue on booting the node. We're working on it.....


-----------------------------------------------------------------------------------

Issue resolved.

Effects - This issue does not effect your relevant products/services
Affecting System - Milano01
Priority - Low
Date - 03/20/2013 12:42 - 03/20/2013 15:22
Last Updated - 03/20/2013 15:22

Network outage (Resolved)

We're aware of a network outage for this server (the whole rack is down). Datacenter has been alerted and we're waiting for a prompt resolution.



----------------------------------------------------


Issue seems resolved.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 03/12/2013 06:05 - 03/12/2013 09:28
Last Updated - 03/12/2013 09:28

Network upgrade (Resolved)

We would like to inform you of a scheduled network maintenance occurring on February 20th, 2013 at 11:00pm EST (Feb 21th at 5am GMT +1). The nature of the maintenance is for network capacity upgrade. The maintenance work will entail adding additional capacity to the DC. The maintenance window will be 30 minutes, but individual customer outages are expected to be less than 15 minutes. Thank you for your understanding and continued business. 


--------------------------------------------------

 Our network maintenance performed last night has been completed. We will have an additional network maintenance scheduled on February 27th, 2013 at 5am GMT+1 to finish the network upgrade. 

Effects - This issue does not effect your relevant products/services
Affecting Other - USA (NC) Services
Priority - Low
Date - 02/27/2013 05:00 - 02/27/2013 14:56
Last Updated - 02/21/2013 17:58

Node issue (Resolved)

There an ongoing issue on this node, techs on site are checking for an hardware issue.


----------------------------------------------------------

It was a ram failure. Ram has been replaced and node is coming back online now.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 02/07/2013 18:08 - 02/07/2013 18:31
Last Updated - 02/07/2013 18:31

DDOS (Resolved)

There's currently an huge ddos attack against the server. We're working to bring it back online.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 01/26/2013 17:24 - 01/26/2013 20:25
Last Updated - 01/28/2013 09:26

Network issue (Resolved)

There is currently an issue with certain routes into the Manchester, UK data center at this time. Our network engineers are currently working to resolve this issue, and restore normal service as soon as possible.

Affected services: vps control panel and UK hosted servers.

Effects - This issue does not effect your relevant products/services
Affecting System - UK (Manchester)
Priority - Critical
Date - 01/25/2013 07:55 - 01/25/2013 10:29
Last Updated - 01/25/2013 09:57

Network issues (Resolved)

There seems to be some net issues in Amsterdam DC, we're investigating.


----------------------------------------

Some IPs are reachable and some are not at this time.

Effects - This issue does not effect your relevant products/services
Affecting System - NL DC
Priority - Critical
Date - 12/14/2012 10:30 - 12/14/2012 15:24
Last Updated - 12/14/2012 12:07

Scheduled maintenance (Resolved)

Scheduled maintenance will be performed on Thursday to Friday night. Services hosted in Amsterdam will go offline for a few minutes.


--------------------------------


Completed.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Low
Date - 11/30/2012 01:45 - 11/30/2012 16:25
Last Updated - 11/30/2012 16:25

Outage (Resolved)

We're aware of an issue on this node. Techs are investigating.


--------------------------


Fixed

Effects - This issue does not effect your relevant products/services
Affecting System - NC01 node (USA)
Priority - High
Date - 11/06/2012 10:05 - 11/06/2012 11:17
Last Updated - 11/06/2012 11:17

NL (Resolved)

There's an ongoing ddos attack. all nl services are down atm 


----------------------------------------------------


Just an update, ddos is stopped, btw we're having issues bringing back online the node Amsterdam01. We're stil working on it


----------------------------------------------------


Still under a few hundreds mbps dos


----------------------------------------------------


Datacenter has decided to take this node/server offline for the next 24 hours because the dos was affecting their whole network. Unfortunately there's nothing we can do right now.


----------------------------------------------------


Node is back online. Services will be reactivated gradually in the next 24 hours


----------------------------------------------------


All services are back online except shared hosting


----------------------------------------------------


The shared/reseller accounts machine that is still down was the target of the ddos. If we bring it back online and ddos happens again, all node customers will go down (DC stated that they'll shut it down for 3 days this time). We're considering moving that machine somewhere else, where it cannot affect others.


----------------------------------------------------


We've decided to put it back online. Fingers crossed. 'till now all fine...


----------------------------------------------------


Everything went fine. 

Effects - This issue does not effect your relevant products/services
Affecting System - Amsterdam01
Priority - Critical
Date - 10/19/2012 20:39 - 10/22/2012 01:24
Last Updated - 10/22/2012 01:22

Networking issues (Resolved)

We're experiencing some network issues with this server. techs are investigating...


---------------------------------------------------------------------

Looks like the server is under an intermittent ddos attack

---------------------------------------------------------------------

Hopefully the attack is under control...

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Medium
Date - 10/12/2012 16:39 - 10/13/2012 03:20
Last Updated - 10/13/2012 03:20

Emails from us (Resolved)

We're testing a new email delivering method.

We've noticed that some emails haven't been sent because of some errors on our smtp server.
We're looking into the issue.

Effects - This issue does not effect your relevant products/services
Affecting System - All emails coming from our billing system
Priority - Low
Date - 10/12/2012 13:45 - 10/14/2012 22:50
Last Updated - 10/12/2012 13:48

Chicago issues (Resolved)

We're aware of networking issues for the Chicago based services.

Techs are working on it...


---------------------------------------------------


We're performing emergency maintenance to fix the issue. There will be a bit of downtime.


---------------------------------------------------

Effects - This issue does not effect your relevant products/services
Affecting Other - Chicago DC
Priority - Critical
Date - 09/29/2012 21:29 - 10/01/2012 08:54
Last Updated - 09/30/2012 23:33

Scheduled maintenance (Resolved)

Maintenance is being performed on this server.

Effects - This issue does not effect your relevant products/services
Affecting System - Vps node FR 01
Priority - Medium
Date - 08/23/2012 17:11 - 08/23/2012 17:45
Last Updated - 08/23/2012 17:28

Nameservers issue on bibi91 (Resolved)

There's currently an ongoing issue with sites using bibi91.info nameservers. 

If you can't see your website please contact support.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 08/11/2012 13:47 - 08/12/2012 12:08
Last Updated - 08/11/2012 13:50

Motherboard replacement (Resolved)

There have been about 2 hours downtime this past night on this server.

An emergency intervention was necessary and its motherboard has been replaced.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 06/27/2012 02:19 - 06/27/2012 04:25
Last Updated - 06/27/2012 11:40

Network issue (Resolved)

Chicago facility is currently experiencing problems with connectivity. We work closely with datacenter staff to resolve the issue and will update you once it is resolved.

 

----------------------------

 

Back online

 

----------------------------

 

In order to address recent network outage Continuum Data Center is performing emergency maintenance and troubleshooting of the problem. Short interruptions in service expected during the day.

We are sorry for the inconveniences and thank you for understanding.

 

----------------------------

 

Again 10:22AM CST we had a number of our edge switches, located at our POP at 350 E. Cermak (Equinix), shut down due to excessive heat (109+ C).

Our Edge switches returned to normal temps at ~11:25 AM CST and traffic began passing normally.

We are continuing to investigate a possible IOS issue that may be causing these issues along with the heat alarms.

Thank you for your patience and we apologize for the inconvenience.

Effects - This issue does not effect your relevant products/services
Affecting Other - Chicago DC
Priority - Critical
Date - 06/05/2012 15:00 - 06/05/2012 22:34
Last Updated - 06/05/2012 19:26

Network outage (Resolved)

There's currently a network issue on Milan (IT) DC

---------------------------------------------------------

Predicted downtime 90 minutes from now, switches have to be replaced.

---------------------------------------------------------

Back online

---------------------------------------------------------

Issues still going on...

---------------------------------------------------------

The new switch that has been installed this morning was defective. The technician who went to take another one from the warehouse is into the Milan's traffic jam caused by the Pope's visit...

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 06/01/2012 10:22 - 06/02/2012 03:40
Last Updated - 06/01/2012 17:55

Net issues (Resolved)

There are some connectivity issues that are being investigated.

Effects - This issue does not effect your relevant products/services
Affecting System - Amsterdam DC
Priority - High
Date - 04/19/2012 18:31 - 04/19/2012 21:33
Last Updated - 04/19/2012 18:31

Issue on server (Resolved)

There's currently an abusive issue on this server.

It'll be back shortly...

---------------------------------------------------------------

It'll be back in a few minutes.

---------------------------------------------------------------

Datacenter is taking forever to reactivate the server, unfortunately we have no control over it and can do nothing other than complain...

---------------------------------------------------------------

Back online

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 04/19/2012 12:06 - 04/19/2012 14:37
Last Updated - 04/19/2012 14:32

Network issue (Resolved)

There seems to be some network issues with the Milan DC.

----------------------------------------------------------------------------

The outage was on the primary router on datacenter. The redundant system failed and this caused about 3 hours of no connectivity.

All systems have been restored. There are still minor connectivity issues with AMS-IX and Atrato that are being resolved.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 04/08/2012 18:58 - 04/19/2012 12:21
Last Updated - 04/09/2012 08:50

Partial network outage (Resolved)

One of our routers in the datacenter is currently causing some instability in the network. While we are reachable from most locations, you could experience some connectivity problems at this moment. Engineers are already in the datacenter looking at this problem.

 

------------------------------------------------------------------------------------------------------------------------------------------

 

The network issue has been fully resolved for a while now (2 hours ago). It was caused by a full routing table on one of the core routers. We will make sure necessary changes will be made to prevent this from happening again. We are sorry for any possible inconvenience.

Affected: 46.182.104.0/21 and 15-25% of its traffic (mainly AMS-IX).
Downtime: 15 minutes and 1.5 hours for AMS-IX traffic.

Effects - This issue does not effect your relevant products/services
Affecting System - Services in Amsterdam
Priority - High
Date - 04/06/2012 15:10 - 04/06/2012 17:28
Last Updated - 04/06/2012 17:28

cPanel and MySql upgrade (Resolved)

cPanel has been upgraded to a newer version. Clients may notice some changes in user interface and functionalities.

----------------------------------------------------------------------------------------------------------------------------------------

MySql is scheduled to be upgraded to version 5.5 in the following days.

Effects - This issue does not effect your relevant products/services
Affecting System - All Shared/Reseller servers
Priority - Low
Date - 03/08/2012 23:21 - 04/01/2012 15:19
Last Updated - 03/08/2012 23:24

Network issue (Resolved)

Since this morning there seems to be some networking issues causing this server to be very slow.

 

-----------------------------------------------------------------------------------------------------------------

 

Issue resolved.

 

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - High
Date - 03/04/2012 13:19 - 03/04/2012 14:13
Last Updated - 03/04/2012 14:13

Server issue (Resolved)

There's currently an outage on this server, it's being investigated...

 

----------------------------------------------------------------------------------

 

Running hardware test...

 

----------------------------------------------------------------------------------

 

Hardware tested fine, further investigating...

 

----------------------------------------------------------------------------------

 

Looks like disks arrays got corrupted. Reinstallation required. Unfortunately this mean this server will have to stay down for a few hours while we reinstall and reconfigure everything. Once this is done we'll have to restore the accounts from last night's backups and (at least for bigger sites, 50+GB, this will take even more).

Smaller sites will be back online in a few hours, bigger sites will be back tomorrow.

If you have an account on this server please open a support ticket for updates.

 

----------------------------------------------------------------------------------

 

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 02/14/2012 16:32 - 02/16/2012 23:30
Last Updated - 02/14/2012 18:44

Node issue (Resolved)

There's an issue on Chicago new node.

It's being investigated...

---------------------------------------------------------------

 

Issue has been resolved.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 02/05/2012 00:48 - 02/05/2012 02:58
Last Updated - 02/05/2012 02:58

Amsterdam migration (Resolved)

All accounts on Amsterdam's Cloud02 are being copied to the new node.

If everything goes fine tomorrow morning accounts will be synced one last time from the "old" node and then the IPs will be switched.

As announced by email, there will be some minutes of downtime.

------------------------------------------------------------------------------

Migration completed, not a single second of downtime :)

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Low
Date - 01/29/2012 17:20 - 01/30/2012 10:08
Last Updated - 01/30/2012 10:08

Outage (Resolved)

There's currently an outage on this server. Updates to follow shortly...

--------------------------------------------------------------------------------

Service is back online

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 01/23/2012 16:09 - 01/23/2012 16:46
Last Updated - 01/23/2012 16:46

Emergency reboot (Resolved)

An emergency reboot was necessary, cloud is booting now and will be back shortly.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 01/16/2012 10:14
Last Updated - 01/23/2012 16:30

Network issued on Italian DC (Resolved)

There's currently a network outage on the Miland DC. It is being fixed...

 

-------------------------------------------------------------------------------------

 

Routers have been rebooted and all servers are back online.

Effects - This issue does not effect your relevant products/services
Affecting Other - Milan DC
Priority - Critical
Date - 01/07/2012 14:06 - 01/07/2012 14:14
Last Updated - 01/07/2012 14:14

cloud04 IT issue (Resolved)

The cloud04 is currently down. The issue is being investigated.

 

------------------------------------------------------------------------------

 

Issue resolved, node is booting.

 

The issue was caused by the driver of the virtual network card that crashed.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 12/07/2011 15:05 - 12/07/2011 15:39
Last Updated - 12/07/2011 15:39

Amsterdam cloud outage (Resolved)

We just had an outage on a part of our Cloud in Amsterdam. The outage was caused by a hickup in the internal network which caused hypervisors to be marked offline incorrectly. We are currently investigation the exact technical details and will make sure this will not happen again.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 12/06/2011 16:03
Last Updated - 12/06/2011 16:04

Issue on Amsterdam cloud (Resolved)

There's currently an issue that is being investigated on the Amsterdam cloud.

 

----------------------------------------------------------------------------------------

 

Issue resolved.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 12/04/2011 17:22 - 12/04/2011 20:14
Last Updated - 12/04/2011 20:14

ddos (Resolved)

Node is under ddos attack

 

-----------------------------------------------------

 

Issue resolved.

 

-----------------------------------------------------

 

Under attack again...

 

-----------------------------------------------------

 

Issue resolved

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 11/13/2011 14:00 - 11/13/2011 20:46
Last Updated - 11/13/2011 20:46

ddos (Resolved)

A ddos attack caused a bit of downtime on this node (about 15 minutes). This has already been addressed.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 11/09/2011 22:12 - 11/09/2011 22:13
Last Updated - 11/09/2011 22:13

Php update (Resolved)

Php is being updated to the latest version (5.3.x)

 

-----------------------------------------------------

 

This has been completed.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Low
Date - 11/02/2011 19:05 - 11/02/2011 19:31
Last Updated - 11/02/2011 19:31

Php update (Resolved)

Php is being updated to the latest version (5.3.x)

 

---------------------------

 

Completed

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Low
Date - 11/02/2011 18:57 - 11/02/2011 19:24
Last Updated - 11/02/2011 19:24

Load issues (Resolved)

There are some load issues on this vm. Investigating...

----------------------------------------------------------------------------------

Some abusive users caused the high load.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 11/01/2011 17:28 - 11/01/2011 17:56
Last Updated - 11/01/2011 17:56

Scheduled downtime (Resolved)

Cloud03 will be down some minutes today for manteinance (we need to reboot it).

 

----------------------------------------------------------------------------------------------------

 

This has been completed. Total downtime about 5 minutes.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - High
Date - 10/30/2011 11:48 - 10/30/2011 12:06
Last Updated - 10/30/2011 12:06

IT cloud offline (Resolved)

There have been some technical issues this night with the Milan cloud.

Everything is back to normality now.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 10/20/2011 08:39
Last Updated - 10/20/2011 08:40

Scheduled reboot (Resolved)

The vserver will be rebooted shortly for mantainance (to increase resources/performance).

 

----------------------------------------------------------------------------------------------------------

 

This has been completed.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - High
Date - 10/10/2011 17:16 - 10/10/2011 17:49
Last Updated - 10/10/2011 17:49

Main BibiHost website mantainance (Resolved)

Our main website is scheduled to be moved to a new server over the weekend (probably saturday or sunday).

Nothing will change for you, it's just our website that may be unavailable or the clients area be closed (under mantainance) while the dns changes propagates.

No clients sites or services will be affected in any way.

 

---------------------------------------------------------------------------------------------------------------

 

Aug 14 - This has been postponed to next week.

 

---------------------------------------------------------------------------------------------------------------

 

Aug 20 - This will be done today/tomorrow...

 

---------------------------------------------------------------------------------------------------------------

 

Aug 21 - This has been completed

Effects - This issue does not effect your relevant products/services
Affecting Other - Main BibiHost website
Priority - Low
Date - 08/12/2011 08:56 - 08/24/2011 14:39
Last Updated - 08/21/2011 12:17

Booting issues on some Win vps (Resolved)

Some clients have reported booting issues on their Windows vps. A lot of Windows vps, on different nodes, are affected by this problem.

Techs are working to fix the issue.

Effects - This issue does not effect your relevant products/services
Affecting System - Windows vps, all locations
Priority - High
Date - 08/17/2011 21:18 - 08/18/2011 07:54
Last Updated - 08/17/2011 21:19

Cloud reboot (Resolved)

The Milan cloud will be rebooted today for a scheduled ram and disk upgrade.

Downtime will be about 5 minutes.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Low
Date - 07/20/2011 08:51 - 07/20/2011 16:58
Last Updated - 07/20/2011 08:54

Power issues (Resolved)

This afternoon, the newest UPS system (installed in 1ST QTR 2011) in our Scranton PA facility malfunctioned. This caused huge spikes throughout the power grid, and led to some loads being dropped in the facility. We are currently running without UPS power in a portion of the facility to bypass the problematic UPS, as we await the arrival of repair technicians to resolve the problem. There may be additional temporary losses of power until the issue is repaired.

Effects - This issue does not effect your relevant products/services
Affecting Other - Scranton
Priority - High
Date - 05/31/2011 22:10 - 06/04/2011 09:09
Last Updated - 05/31/2011 22:10

Network outage UK dc (Resolved)

There seems to be a network outage on the Manchester DC. Techs are working on it...

----------------------------------------------------------------------------------------

Everything should be back to normality in minutes....

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 05/15/2011 22:19 - 05/17/2011 19:27
Last Updated - 05/15/2011 22:44

Network issue (Resolved)

The networking department did identify an issue with one of our uplink providers and has contacted them in regards to it...

 

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Medium
Date - 05/07/2011 17:12 - 05/08/2011 04:16
Last Updated - 05/07/2011 17:13

Amsterdam cloud issue (Resolved)

A failure happened on the Amsterdam cloud that caused about 1 hour downtime. The exact details of the failure will follow...

 

------------------------------------------------------------------------------------------------------------------------

 

One hypervisor had SAN connectivity issues. We have migrated these VMs to different hypervisors.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 05/06/2011 17:53 - 05/07/2011 00:50
Last Updated - 05/06/2011 18:00

Varnish removed + Apache rebuild (Resolved)

Varnish cache has been removed from this server.

Apache is being rebuilt.

No downtime is expected.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Medium
Date - 05/04/2011 09:03 - 05/04/2011 10:06
Last Updated - 05/04/2011 09:04

Scheduled maintenance on SAN3 (Resolved)

The Chicago cloud is undergoing scheduled mantainance. It'll be back shortly...

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - High
Date - 04/27/2011 07:20 - 04/27/2011 07:44
Last Updated - 04/27/2011 07:21

Unplanned downtime, resolved (Resolved)

The Chicago Cloud has been unreachable today.

Nothing weird, simply no techs have been alerted because of a malfuncion on the monitoring system.

We're extremely sorry about this, unfortunately this is one of those things that, as much as you try to prevent, just happen.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 04/25/2011 23:12
Last Updated - 04/25/2011 23:17

Vserver rebuilt (Resolved)

Vserver is being rebuilt with higher resources. It'll be down during the process...

 

-----------------------------------------------------------------------------------------------------

 

There's a little issue rebilding the machine, btw it should be up shortly....

 

-----------------------------------------------------------------------------------------------------

 

Unfortunately, because of rebuilding issues it took forever. Btw, at least, it's back online now.

We're going to further investigate on this issue, normally a cloud upgrade should take 1 to 2 minutes (the time to reboot + a bit to add resources)!

We're extremely sorry for the inconvenience :(

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - High
Date - 04/22/2011 23:49 - 04/23/2011 01:52
Last Updated - 04/23/2011 01:52

Network issue (Resolved)

We're currently experiencing a network issue on Scranton DC...

Effects - This issue does not effect your relevant products/services
Affecting System - Scranton DC
Priority - High
Date - 04/08/2011 07:41 - 04/08/2011 09:37
Last Updated - 04/08/2011 07:42

Chicago Cloud - SAN mantainance (Resolved)

We are moving everyone off of SAN5 because of issues with hardware. We have brought up another SAN that addresses and fixes these issues and attached it to the same Pool that the VMs are currently running on. We will be moving each VM disk to the new SAN one by one to ensure that its done properly. It will not be done in any particular order. The moves are starting this afternoon (around 4PM) and will continue until everyone is migrated. Each migration will take around 30 minutes to 1 hour.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - High
Date - 03/30/2011 21:55 - 04/06/2011 15:53
Last Updated - 03/30/2011 21:59

IP block unreachable from some networks (Resolved)

There seems to be some problems reaching this server with IPs from this range 46.37.180.xxx

Server is online and working fine with other IP ranges (eg 178.238.131.xxx).

Technicians are investigating the issue...

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - High
Date - 03/29/2011 02:51
Last Updated - 03/29/2011 05:42

Scheduled upgrade (Resolved)

A scheduled upgrade is being performed on this virtual server cloud based. A reboot is necessary so it'll be down for a few minutes...

 

---------------------------------------------------------------------------

 

This has been completed.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - High
Date - 03/28/2011 16:22 - 03/28/2011 16:57
Last Updated - 03/28/2011 16:57

Server unreachable (Resolved)

Server is unreachable. Techs are investigating...

----------------------------------------------------------------------

Problem solved. Server is booting...

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 03/17/2011 19:24 - 03/17/2011 19:51
Last Updated - 03/17/2011 19:51

Main BibiHost.com site has been moved (Resolved)

Our main site has been moved into the cloud.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Low
Date - 03/15/2011 13:55
Last Updated - 03/15/2011 13:57

Shared accounts down (Resolved)

An issue is going on with the shared hosting cloud based.

Cloud is up, btw the virtual server hosting shared accounts has some problems that are being investigated...

--------------------------------------------------------------------------------------------------------

The virtual server has had to be rebuilt. Accounts have been restored (unfortunately not all) and everything is back to normality.

This all took a few hours because of a bad sequence of coincidences and we are extremely sorry for the downtime and any problem it may have caused.

 

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 03/11/2011 17:02 - 03/12/2011 14:26
Last Updated - 03/11/2011 19:38

Network problems on USA and UK servers (Resolved)

Same clients have reported difficulties to reach their websites from some locations. Our tests have confirmed this issue affects most of our UK and USA servers.

It doesn't seems to be a server related problem, most likely a network outage on some routes.

Engineers are investigating...

This issue affects all services hosted on UK and USA servers so shared hosting, vps and dedicated servers...

-----------------------------------------------------------------------------------------------------------------------------

We are seeing issues with routes for a few of our clients.
Some of them have been related to a routing issue out of our network influence. 
Each network issue must be handled on a case by case basis so it may take some time...

Effects - This issue does not effect your relevant products/services
Affecting Other - Most USA and UK servers
Priority - High
Date - 03/06/2011 11:16 - 03/09/2011 21:26
Last Updated - 03/07/2011 06:02

Testing... :) (Resolved)

Hello {$client_name}

with the porpouse of boosting sites performance and decrease the average load of the server, we've implemented a module called Varnish into the bibi89 server.

This has been installed on this server only as we'd like to give it a try before putting it in production on all servers.

What is it and how does it work?

Varnish stores web pages in memory so that web server do not have to create the same web page over and over again. The web server only recreates a page when it is changed. Varnish can serve web pages much faster than any application server; giving the website a significant speed up.

 

We've checked a few websites hosted on this server and all seems to work fine. Also, server's load has decreased and pages load have dramatically improved, expecially on busy sites.

We'd like to receive a feedback from you, if you noticed improvements on your site's load time or any error.

 

Thank you :)

BibiHost Team

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Low
Date - 03/03/2011 21:39 - 03/05/2011 10:25
Last Updated - 03/03/2011 21:40

Mantainance today (Resolved)

We've scheduled mantainance on this server today. Nothing weird or that may cause downtime, anyway there's always a minimum risk of minutes of server unavailability.

We've also scheduled a surprise that should boost sites performances and reduce server's load drastically but I don't know if this will be implemented today ;)

Updates will follow.....

------------------------------------------------------------------------------------------------------------------

This has been completed. The "surprise" has been implemented too :)

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Low
Date - 03/03/2011 18:55 - 03/03/2011 21:08
Last Updated - 03/03/2011 21:08

Major server issues, accounts will be moved (Resolved)

There're major load issues with this server that can't be resolved. For this reason all accounts will be moved to another server.

------------------------------------------------------------------------------------------------------------------------

 

All sites have been moved from bibi88 to bibi89 server.

Please check your email for details

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 03/01/2011 12:26 - 03/03/2011 19:01
Last Updated - 03/01/2011 21:14

500 internal server error (Resolved)

Apache is being recompiled. Fcgi instead of suphp as php handler as a recent speed improvement caused this memory issue.

All sites should be back shortly.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 03/01/2011 16:05 - 03/01/2011 16:19
Last Updated - 03/01/2011 16:07

Apache recompile (Resolved)

Apache will be recompiled on this server to enhance performance.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Medium
Date - 02/28/2011 19:24 - 02/28/2011 20:35
Last Updated - 02/28/2011 19:25

Server problems (Resolved)

A recent update caused problems with apache. It is being fixed...

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 02/27/2011 16:11 - 02/27/2011 16:38
Last Updated - 02/27/2011 16:12

Urgent mantainance (Resolved)

Urgent mantainance is being performed on this server. A bit of downtime is expected...

------------------------------------------------------------------------------------------------------

24 Feb update: We've tweaked, optimized and monitored the server for 24 hours. The load problem affecting this server lately seems to be an IO related issue. DC technicians are investigating the issue and should come up with a solution shortly...

------------------------------------------------------------------------------------------------------

26 Feb update: Server runs decently but it's not what we want, we want it fast! Still working on it today...

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 02/23/2011 18:33 - 02/28/2011 05:55
Last Updated - 02/26/2011 15:07

bibi33 down (Resolved)

 

The node is down for emergency maintenance. Our engineering team is working hard to get the node back online as soon as possible. 
-------------------------------------------------------------------------------------------

This has been completed.

 

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 02/16/2011 20:21 - 02/16/2011 20:50
Last Updated - 02/16/2011 20:50

Server mantainance (Resolved)

This server is undergoing scheduled mantainance. No downtime is planned.

-------------------------------------------------------------------------------------

Feb 16: Planned mantainance this evening too.

Update: MySQL has been upgraded. We're now recompiling apache to work with the new version, it should take about 10 minutes. In the meantime you may receive "500 error".

Update: This has been completed

-------------------------------------------------------------------------------------

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Medium
Date - 02/15/2011 19:06 - 02/16/2011 19:04
Last Updated - 02/16/2011 19:04

Temporary problem (Resolved)

Apache is being recompiled because of an error. Sites may temporary not work.

Everything should be back to normality shortly...

-------------------------------------------------------------------------------------------

Everything is back to normality now.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 02/14/2011 20:58 - 02/14/2011 23:19
Last Updated - 02/14/2011 23:19

bibi33 down (Resolved)

There's currently an issue with bibi33 server. We're working to have it back up running asap.

 

---------------------------------------------------------------------------------------------------------

 

A full backup from mid January has been rstored. Now it's working fine. We're now restoring each user backup from yesterday's backups. Everything should return to normality shortly.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 02/05/2011 05:02 - 02/08/2011 19:14
Last Updated - 02/05/2011 05:49

Server issue (Resolved)

Node bibi33 (LA) is down, techs are investigating...

----------------------------------------------------------------------

Problem fixed.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 01/04/2011 13:35 - 01/04/2011 13:59
Last Updated - 01/04/2011 13:59

Mantainance (Resolved)

This server is going under mantainance today and will probably stay offline a bit while engineers are working on it.

We appreciate your patience and understanding through this process.

 

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - High
Date - 12/28/2010 08:42 - 12/28/2010 17:47
Last Updated - 12/28/2010 08:49

Server outage (Resolved)

There seems to be an outage on this node. Techs are investigating...

 

----------------------------------------------------------------------------

 

Node is back online.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 11/11/2010 21:36 - 11/12/2010 00:03
Last Updated - 11/11/2010 22:40

Server outage (Resolved)

There seems to be an outage on this node. Techs are investigating...

 

----------------------------------------------------------------------------

 

Node is back online.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 11/11/2010 21:37 - 11/12/2010 00:03
Last Updated - 11/11/2010 22:40

ffmpeg mantainance (Resolved)

We're doing some ffmpeg and related modules mantainance....

FFmpeg may not function properly.

 

-----------------------------------------------------------------------------

 

Still working of this....

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Medium
Date - 10/18/2010 21:42 - 10/24/2010 19:56
Last Updated - 10/21/2010 06:00

Server upgrade (Resolved)

Server is going to be upgraded from 100Mbps to 1Gbps during this weekend :)

A bit of downtime is scheduled for the hardware upgrade...

 

--------------------------------------------------------------------

 

This upgrade will be performed on monday...

 

--------------------------------------------------------------------

 

Done :D

 

# wget http://cachefly.cachefly.net/100mb.test
--2010-10-19 14:47:43-- http://cachefly.cachefly.net/100mb.test
Resolving cachefly.cachefly.net... 205.234.175.175
Connecting to cachefly.cachefly.net|205.234.175.175|:80... connected.
HTTP request sent, awaiting response... 200 OK
Length: 104857600 (100M) [application/octet-stream]
Saving to: `100mb.test'

100%[===================================================================================================================>] 104,857,600 69.0M/s in 1.4s

2010-10-19 14:47:44 (69.0 MB/s) - `100mb.test' saved [104857600/104857600]

 

------------------------------------------------------------------------------------------------------

 

Just finished doing the last part of mantainance that caused some minute of downtime.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Low
Date - 10/15/2010 23:42 - 10/21/2010 05:59
Last Updated - 10/20/2010 19:41

Server upgrade (Resolved)

Server is going to be upgraded from 100Mbps to 1Gbps during this weekend :)

A bit of downtime is scheduled for the hardware upgrade...

 

--------------------------------------------------------------------

 

This upgrade will be performed on monday...

 

--------------------------------------------------------------------

 

Done :D

 

# wget http://cachefly.cachefly.net/100mb.test
--2010-10-19 14:47:43-- http://cachefly.cachefly.net/100mb.test
Resolving cachefly.cachefly.net... 205.234.175.175
Connecting to cachefly.cachefly.net|205.234.175.175|:80... connected.
HTTP request sent, awaiting response... 200 OK
Length: 104857600 (100M) [application/octet-stream]
Saving to: `100mb.test'

100%[===================================================================================================================>] 104,857,600 69.0M/s in 1.4s

2010-10-19 14:47:44 (69.0 MB/s) - `100mb.test' saved [104857600/104857600]

 

------------------------------------------------------------------------------------------------------

 

Just finished doing the last part of mantainance that caused some minute of downtime.

 

 

 

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Low
Date - 10/15/2010 23:40 - 10/21/2010 05:59
Last Updated - 10/20/2010 19:41

Main BibiHost website move (Resolved)

Our main website (bibihost.com) will be moved on a different server today so it may appear down while nameservers propagates...

Also, billing area will be turned off temporarely.

-----------------------------------------------------------------------------------

 

This has been completed :)

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Medium
Date - 10/17/2010 10:58 - 10/17/2010 20:31
Last Updated - 10/17/2010 16:56

New OS images for vps (Resolved)

We're rebuilding the list of operating systems that can be installed on our vps node.

CentOS 5 will be updated to CentOS 5.5

Fedora 12 updated to Fedora 13

addition of Ubuntu 10.04 and 10.04 minimal

CentOS 5 + cPanel updated to CentOS 5.5 + latest version of cPanel

CentOS 5 + Webmin updated to CentOS 5.5 + Webmin

This will be completed this weekend :)

 

-------------------------------------------------------------------------------

 

This has been completed :)

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Low
Date - 10/16/2010 18:52 - 10/17/2010 20:31
Last Updated - 10/17/2010 16:44

Server move (Resolved)

We've started moving sites to the new server....

 

-----------------------------------------------------------------------------------

 

This has been completed. Old server bibi77 will be dismissed on monday.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Medium
Date - 10/15/2010 23:39 - 10/19/2010 11:46
Last Updated - 10/16/2010 13:52

Server down (Resolved)

Server is down.....

---------------------------

Back online :)

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 08/19/2010 17:44 - 08/19/2010 19:26
Last Updated - 08/19/2010 19:26

Server reboot (Resolved)

A server reboot is scheduled today.

--------------------------------------------

Done, total downtime 90 seconds :)

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Low
Date - 08/19/2010 15:11 - 08/19/2010 17:14
Last Updated - 08/19/2010 17:14

Server problem (Resolved)

There's something wrong with this server (it keeps rebooting after last update).

Techs are investigating....

----------------------------------------------------------------------------------------------

Problem fixed.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 08/18/2010 14:03 - 08/18/2010 16:34
Last Updated - 08/18/2010 16:34

S.M.A.R.T Errors on /dev/sda (Resolved)

Errors on sda of this server have been reported by the system. This server runs on a raid1 setup so there shouldn't be problems, anyway we're investigating the issue and it may be necessary to change a drive.

 

-----------------------------------------------------------------------------------------------------

We're keeping it monitored.

 

-----------------------------------------------------------------------------------------------------

 

An hardware test is scheduled. We'll have to take the server offline for about 1 hour.

 

-----------------------------------------------------------------------------------------------------

 

Tests in progress...

 

-----------------------------------------------------------------------------------------------------

 

 Tests terminated, everything looks fine. Hopefully it was only a false alarm :)

 

-----------------------------------------------------------------------------------------------------

Update august 10: the system keeps logging errors for the drive a of the server. It may be necessary to substitute it and rebuild the raid from drive b today...

 

-----------------------------------------------------------------------------------------------------

Server offline for mantainance

 

-----------------------------------------------------------------------------------------------------

The faulty drive (drive A) has been substituted. We're now rebuilding the new drive (A) from the raid1 drive (drive B)

 

-----------------------------------------------------------------------------------------------------

Drive is rebuilding, the whole procedure should finish in about 90 minutes

----------------------------------------------------------------------------------------------------

Server should be back in about 20 minutes...

----------------------------------------------------------------------------------------------------

Still working on it....

----------------------------------------------------------------------------------------------------

There seems to be major problems on this server. We're restoring the accounts on another server. All of them if not already online, should come up shortly (propagation of nameservers change)

----------------------------------------------------------------------------------------------------

The server will probably be reinstalled tonight/tomorrow morning. We're currently backing up some large account from it that haven't been moved to the temp server. Hopefully we'll be able to put the sites back to it by tomorrow afternoon.

----------------------------------------------------------------------------------------------------

august 13

Reinstalling, configuring, testing the server... Sites will be moved back hopefully later today, most likely tomorrow...

----------------------------------------------------------------------------------------------------

august 14

Accounts are being transferred. Process will finish tonight/tomorrow...

----------------------------------------------------------------------------------------------------

Finished earlier than expected :)

All sites have been moved back to the bibi77 server. Nameservers have been changed and are now propagating.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - High
Date - 07/21/2010 06:07 - 08/18/2010 14:03
Last Updated - 08/14/2010 18:09

Apache malfunctions (Resolved)

Apache is being rebuilt and updated. We're aware that sites in this server aren't working properly. Everything should be back to normality shortly...

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - High
Date - 08/03/2010 09:57 - 08/03/2010 11:28
Last Updated - 08/03/2010 09:59

Work tonight (Resolved)

There will be works on the fiber optic cables between
Roubaix and Brussels, this will cause cuts in
Roubaix/Amsterdam and Roubaix/Frankfurt. The traffic will be rerouted
automatically via Roubaix/London/Amsterdam and Roubaix/Paris/Frankfurt.

Effects - This issue does not effect your relevant products/services
Affecting Other - France DC (France shared and reseller accounts, France dedicated servers)
Priority - Low
Date - 07/27/2010 20:51 - 08/10/2010 06:51
Last Updated - 07/27/2010 20:54

Network problems (Resolved)

We're aware of some network issues on France DC and are investigating....

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 07/12/2010 06:09 - 07/12/2010 06:23
Last Updated - 07/12/2010 06:11

Server outage (Resolved)

A major outage occured. It has been resolved but unfortunately caused an important downtime. We're extremely sorry for this, hopefully it shouldn't happen again.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 05/29/2010 00:29 - 05/29/2010 00:29
Last Updated - 05/29/2010 00:29

Server outage (Resolved)

 

A major outage occured. It has been resolved but unfortunately caused an important downtime. We're extremely sorry for this, hopefully it shouldn't happen again.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 05/29/2010 00:02 - 05/29/2010 00:05
Last Updated - 05/29/2010 00:05

FFmpeg mantainance (Resolved)

 

FFmpeg is going to be recompiled with some more functions and may not work properly in the following hours. This should be completed in 24 hours...

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Low
Date - 05/13/2010 22:07 - 05/14/2010 00:00
Last Updated - 05/17/2010 23:18

ddos attacks (Resolved)

Unfortunately there're always bored people that has nothing better to do than ddosing our bibi77 server causing downtime. Today, even with various protections on the server, it has been ddosed for 1 hour causing downtime to us and all clients sharing this server.

Since we host our main site bibihost.com on this server, we're going to move it to its own dedicated server.

This will isolate it from a server where we host clients meaning if someone attacks us your site will not be affected and viceversa. This will, for sure, improve the uptime (the server where we host our main bibihost website is the only one that has downtime problems. All of our other servers are always up and stable)

Also, if you get banned by the server's firewall for some reason, you'll still be able to connect to our main site to open a support ticket.

We're most likely going to move it during the weekend.

Team @ BibiHost

 

 

-------------------------------------------------------------------------------------------------------------------------------------------------------------------

 

 

This has been completed.

 

 

 

--------------------------------------------------------------------------------------------------------------------------------------------------------------------

 

 

We've also done a step forward in preventing ddos attacks on each of our servers doing a specific tweak of each firewall ;)

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - High
Date - 04/10/2010 01:00
Last Updated - 04/11/2010 03:39

Billing system undergoing mantainance (Resolved)

Our billing system is being updated and may not work properly during the process.

 

---------------------------------------------------------------------------------------------

 

Update finished.

Effects - This issue does not effect your relevant products/services
Affecting Other - Billing
Priority - Low
Date - 04/05/2010 14:23
Last Updated - 04/06/2010 06:45

Emergency reboot (Resolved)

An emergency reboot has been performed this night on our main server.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Critical
Date - 03/22/2010 06:42
Last Updated - 03/22/2010 06:43

Some network problems on USA DC (Resolved)

There seems to be some random network outages today on the USA datacenter. Techs on site are working to solve the issue. Affected clients are on server bibi44, USA dedicated servers and Vps.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - High
Date - 03/17/2010 13:51 - 03/17/2010 00:00
Last Updated - 03/17/2010 21:38

Apache problem (Resolved)

There's currently an issue that causes Apache to fail quite often. This causes random downtimes of a few minutes here and there.

Techs are working to fix it.

 

--------------------------------------------------------------------------------------------------------------------------------------------------

 

Apache has been tweaked, recompiled and updated. Hopefully there shouldn't be problems anymore...

 

---------------------------------------------------------------------------------------------------------------------------------------------------

 

Still working on it...

 

---------------------------------------------------------------------------------------------------------------------------------------------------

 

Still working on it... We already have another server ready in case the issue continues. We'll see in the following hours...

 

---------------------------------------------------------------------------------------------------------------------------------------------------

 

Problem fixed, let's hope ;)

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - High
Date - 03/02/2010 13:40 - 03/03/2010 14:32
Last Updated - 03/03/2010 14:32

Server reboot (Resolved)

Server has been rebooted for mantainance. There have been few minutes of downtime, that's normal.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Medium
Date - 02/03/2010 21:23 - 02/03/2010 00:00
Last Updated - 02/03/2010 21:25

Update of the IOS on the router (Resolved)

Router that serves this server has been updated during the night. The server has been down for 15 minutes.

Effects - This issue does not effect your relevant products/services
Affecting Server -
Priority - Low
Date - 01/19/2010 00:00 - 01/21/2010 06:58
Last Updated - 01/21/2010 06:56

 

rss View RSS Feed

Our official partners